OpenStudy (anonymous):

3x-12y=-60 what is the slope draw it please.

OpenStudy (mathstudent55):

Start by solving for y.

OpenStudy (johnweldon1993):

We can change this into slope intercept form so we can see exactly what the slope is So we want to make THAT look like: \[\large y = mx + b\] because here, we see that 'm' is the slope of the line...and 'b' is the y-interept...so 3x - 12y = -60 if we subtract 3x from both sides of the equation we have -12y = -3x - 60 Now if we divide EVERYTHING by 12...we get \[\large y = \frac{3}{12}x + \frac{60}{12}\] or \[\large y= \frac{1}{4}x + 5\] so now we see that 1/4 is the slope...and 5 is the y-intercept